OpenStudy (anonymous):

in a triangle abc, a and b are the same measure, c is 57 less than either of them, what are the measures. I need assistance on how to do the equation.

OpenStudy (anonymous):

Since you are told that a and b are the same measure, you can say that a=b. You are told that c is 57 degrees less than either a or b; since a=b, you can say c = a-57. Remember that the sum of the angles of a triangle = 180. Now you have the following equation: 180 = a + a + (a-57) 180 = 3a-57 237 = 3a a = b = 79 180-(2*79) = 180-158 = 22 a = 79, b = 79, c = 22

OpenStudy (anonymous):

none of that made any since. but thanks for trying to help